# Generating Model Parameters for Controlling: Parameter Diffusion for Controllable Multi-Task Recommendation

### v1

@@ 6 Related Work
Figure 5: Performances of different conditioning strategies on MovieLens 1M using SASRec as backbone. The results of the “Retrain” algorithm are used as a reference.
Diffusion models. Diffusion probabilistic models (Ho et al., 2020; Song et al., 2020; Nichol & Dhariwal, 2021)have not only achieved significant success in the field of image generation but have also found wide applications in various other areas in recent years, such as video generation  (Ho et al., 2022b), text generation (Li et al., 2022; Gong et al., 2022), etc. Moreover, diffusion models have shown the ability to generate high-quality neural network parameters, achieving comparable or even superior performance to traditionally trained models (Yuan et al., 2024; Wang et al., 2024). These models have also been applied to enhance the accuracy of recommender systems by addressing challenges such as noisy interactions and temporal shifts in user preferences (Wang et al., 2023). In our work, we utilize diffusion models to generate parameters for controllable multi-task recommender systems.
Multi-task learning (MTL) aims to develop unified models that tackle multiple learning tasks simultaneously while facilitating information sharing (Zhang & Yang, 2021; Ruder, 2017). Recent advancements in MTL include deep networks with various parameter sharing mechanisms (Misra et al., 2016; Long et al., 2017; Yang & Hospedales, 2016) and approaches treating MTL as a multi-objective optimization problem (Lin et al., 2019; Mahapatra & Rajan, 2020; Xie et al., 2021). These latter methods focus on identifying Pareto-efficient solutions across tasks, with significant applications in recommender systems (Jannach, 2022; Li et al., 2020a; Zheng & Wang, 2022) Researchers have explored different strategies, from alternating optimization of joint loss and individual task weights to framing the process as a reinforcement learning problem (Xie et al., 2021). The emphasis has shifted from optimizing specific preference weights to finding weights that achieve Pareto efficiency across objectives (Sener & Koltun, 2018; Lin et al., 2019). Some methods utilize attention mechanisms to dynamically allocate computational resources among tasks (Liu et al., 2019). More recent approaches, such as the CMR (Chen et al., 2023), utilize hypernetworks to learn the entire trade-off curve for MTL problems. However, existing works lack of the attention on test-time controllability when receiveing new task requirements. Inspired by the framework of controllable learning (Shen et al., 2024), our novel approach diverges from these existing methods by employing diffusion models to control model parameters at test time, potentially offering greater flexibility and adaptability in handling multi-task learning problems.
7 Conclusions
@@
This paper addresses the critical challenge of adapting recommendation models to dynamic task requirements in real-world applications, where frequent retraining is impractical due to high computational costs. To tackle this problem, we propose PaDiRec, a novel controllable learning approach that enables efficient adaptation of model parameters without retraining by utilizing a diffusion model as a parameter generator. Our approach is model-agnostic, allowing it to integrate with existing recommendation models and enhance their controllability. PaDiRec provides a practical solution for real-time, customizable recommendations in achieving efficient, test-time adaptation.
Appendix A Appendix

### v2

@@ 5.3.1. RQ1: What are the advantages of Diffusion over Hypernetwork in parameter generation?
Figure 8. Performances of different conditioning strategies on MovieLens 1M using SASRec as backbone. The results of the “Retrain” algorithm are used as a reference.
Paragon is designed to adaptively adjust model parameters in an online environment without retraining, enabling it to quickly respond to new task requirements. This places a strong emphasis on the model’s response time. We compared the response times of Paragon and “Retrain” across various backbones and datasets, with the results shown in Table 3. As observed, in all experiments using three different backbones across three datasets, Paragon’s response time was significantly faster than that of “Retrain”. Notably, the response time of “Retrain” correlated with the size of the dataset, whereas Paragon exhibited minimal variation across different datasets. This highlights Paragon’s data-agnostic nature indicating its potential for handling large-scale datasets efficiently.
5.3.3. RQ3: What about the scalability to more utilities such as fairness.
We examined the performance of Paragon under various metric controls, including the widely used fairness metric in recommendation systems, where fairness is defined as the absolute difference (AD) of NDCG@10 between male and female groups. We set the value of fairness dimension in preference weight to $\{0,1\}$, corresponding to unfair and fair settings. As shown in Figure 6 (a), under different levels of control over accuracy and diversity, there is a clear and consistent difference in the models’ AD. when set control signal to fair versus unfair. A smaller AD for the fair setting indicates better inter-group fairness. Meanwhile, as shown in Figure 6 (b), we divided the control signals into two groups: fair and unfair. Within each group, accuracy and diversity remain stably negatively correlated, while the differences in accuracy and diversity between the groups are minimal. This experiment demonstrates that Paragon can effectively achieve control over multiple objectives.
5.3.4. RQ4: How do different conditioning strategies impact the model’s performance?
We investigate the influence of different conditioning strategies aimed at improving the integration of conditions into the denoising model. The detailed structure is shown in Figure 7. Each strategy emphasizes different performance dimensions as depicted in Figure 8. In terms of Hypervolume, all five strategies outperform the “Retrain” approach, with the “Pre&Post” strategy achieving the best results. For Pearson r-a and Pearson r-d, the “Adap-norm” strategy demonstrates the best overall performance, indicating strong consistency with the “Retrain” approach, i.e., high controllability. Additionally, the Hypervolume remains within an acceptable range, suggesting that adding conditions aggregated by an attention mechanism to the layer norm is a promising approach for controllability.
6. Case Study
Table 4. The top-10 recommendation lists under accuracy weights (abbreviated as Acc.) of 0.1 and 0.9. The item order in the table reflects the order in recommendation list. This experiment is conducted on the MovieLens 1M dataset, utilizing SASRec as the backbone.
Acc. Category Item Is Target Item
0.1 Animation, Children’s, Comedy, Musical, Romance Little Mermaid No
0.1 Action, Comedy, Crime, Horror, Thriller From Dusk Till Dawn No
0.1 Adventure, Fantasy, Sci-Fi Time Bandits No
0.1 Animation, Children’s Sword in the Stone No
0.1 Action, Romance, Thriller Desperado No
0.1 Adventure, Children’s, Fantasy Santa Claus No
0.1 Horror, Sci-Fi Invasion of the Body Snatchers No
0.1 Film-Noir, Mystery,Thriller Palmetto No
0.1 Action, Comedy Twin Dragons No
0.1 Film-Noir Sunset Blvd. No
0.9 Horror Birds Yes
0.9 Drama Cider House Rules No
0.9 Comedy, Romance Annie Hall No
0.9 Action, Comedy, Crime, Horror, Thriller From Dusk Till Dawn No
0.9 Drama, Romance Girl on the Bridge No
0.9 Animation, Children’s, Comedy, Musical, Romance Little Mermaid No
0.9 Comedy Road Trip No
0.9 Comedy, Drama Chuck & Buck No
0.9 Horror, Sci-Fi Invasion of the Body Snatchers No
0.9 Animation, Children’s Sword in the Stone No
To better illustrate the utility of Paragon, we present the top-10 recommendation lists under two sets of preference weights. We compared the top-10 recommendation lists between an accuracy weight of 0.1(diversity weight is 0.9) and an accuracy weight of 0.9(diversity weight is 0.1). Notably, when the accuracy weight is 0.1 (indicating a high preference for diversity), items covering more categories are ranked higher, but the list does not include the target item, indicating poor accuracy. Conversely, with an accuracy weight of 0.9, the target item is ranked in the top 1 position within the recommendation list, but the top items cover fewer categories.
7. Conclusions
@@
This paper proposed Paragon to address the critical challenge of adapting recommendation models to dynamic task requirements in real-world applications, where frequent retraining is impractical due to high computational costs. As a novel controllable learning approach, Paragon conditionally generate parameters instead of retraining. Overall, Paragon provides a practical solution for real-time, customizable recommendations, which provide a feasible approach for controllable learning.
Acknowledgements.
This work was partially supported by the National Natural Science Foundation of China (No. 62376275, 62472426). Work partially done at Beijing Key Laboratory of Research on Large Models and Intelligent Governance, and Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE. Supported by fund for building world-class universities (disciplines) of Renmin University of China. Supported by the Research Funds of Renmin University of China (RUC24QSDL016).